OpenStudy (anonymous):

Solve for m: w = m - 5L. A. m = w - 5L B. m = w + 5L C. m = (w - 5)/L D. m = w/L + 5

jimthompson5910 (jim_thompson5910):

w = m - 5L w+5L = m - 5L+5L w + 5L = m m = w + 5L